OpenStudy (anonymous):

Finding Sin(55°) I found that Sin(55°) = Sin(11Pi /36) ? Help please, without using the calculator!!

OpenStudy (ybarrap):

you can try this you know that sin 45 is sqrt(2)/2 and sin(60) is sqrt(3)/2 sin(55) is between them The closer to 55 the better so, we weight by the inverse of the distance between these angles we'll give the sin 45 a weight of 1/(55-45) = 1/10 and sin 60 a weight of 1/(60-55) = 1/5 -- this is larger because 60 is closer to 55 than 45. So, cos(55) is about ((1/10)*sin 45 + (1/5)*sin(60)) / ( 1/10 + 1/5) Using this technique will get you to the exact value of 1% (without a calculator - or calculus). If you know calculus, you can use the taylor expansion of cos(x), which will be as accurate as you want: sin(x) = x-x^3/6+x^5/120-x^7/5040+x^9/362880-x^11/39916800+O(x^12)

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (ybarrap):

NP A couple of typos - *So, sin(55) is about ((1/10)*sin 45 + (1/5)*sin(60)) / ( 1/10 + 1/5) *If you know calculus, you can use the taylor expansion of sin(x), which will be as accurate as you want: sin(x) = x-x^3/6+x^5/120-x^7/5040+x^9/362880-x^11/39916800+O(x^12)